# Feedback Optimization with State Constraints through Control Barrier Functions

### v1

@@ I Introduction
  • We prove local exponential stability of local optima, assuming that the optimum lies in the interior of the state constraint set;
@@
The existing literature on feedback optimization either disregards state constraints, enforces them only asymptotically, e.g. [4, 10, 11, 12, 13], or considers only static plants, e.g. [14, 15]. In the latter two cases, which are equivalent, state constraints can be enforced through input constraints, as there is a one-to-one correspondence between state and input. Applying these methods to dynamic plants may lead to constraint violations in the transients. In contrast, the method proposed here is the first one to enforce state constraints at all times for dynamic plants. Nonetheless, our method comes with local stability guarantees, compared to other works (e.g., [9]) that guarantee global convergence to critical points or optimizers. Whether our method can ensure global convergence remains a topic of future research.
SGFs were first introduced in [16], as dynamical systems that solve constrained optimization problems. These are modified gradient flows, that employ CBFs, to enforce constraint satisfaction. Further, in [19, 20], it was shown how SGFs are continuous approximations of projected gradient flows. SGFs have been used in reinforcement-learning to ensure safety in [21, 22]. In the context of feedback optimization, they have been used as controllers interconnected with dynamic plants in [15, 13], where however state constraints are only asymptotically enforced. Our work is the first to combine SGFs with high-order CBFs. In fact, high-order CBFs are the key to enforce state constraints, since these are of relative degree greater than one with respect to the controller dynamics and cannot be handled by standard SGFs interconnected to a plant, as SGFs (and projected gradient flows) manipulate only the controller’s vector-field. High-order CBFs enable modifications of the controller’s vector field that enforce state constraints.
Finally, while we focus on feedback optimization, there are other methods targeting similar problems, such as extremum seeking (ES; e.g. [23, 24]) and online convex optimization (OCO; e.g. [25, 26, 27]). Like existing feedback-optimization methods, ES methods that address state constraints either cast them as input constraints, implicitly assuming that the plant operates at the steady-state, or consider static plants (which are equivalent approaches); see [23, 24] and references therein. For differences between feedback optimization and ES the reader is referred to [1]. With respect to OCO, the main differences are that OCO generally considers discrete-time plants and focuses on finite-horizon regret guarantees, rather than asymptotic convergence to the optimizer.
@@ II Preliminaries
For $r\in\mathbb{Z}_{>0}$, we denote $[r]:=\{1,\ldots,r\}$. We adopt the following conventions on dimensions: given $g:^{n}\times^{m}\to^{k}$, $(x,u)\mapsto g(x,u)$, continuously differentiable, then $\dfrac{\partial g}{\partial x}(x,u)\in^{k\times n}$; given $g:^{n}\to$, $x\mapsto g(x)$, continuously differentiable, then $\nabla g(x)=(\frac{\partial g}{\partial x}(x))^{\top}\in^{n}$. Consider the parametric optimization problem
@@
where $\Phi,a_{i}:^{n}\times^{p}\to$ and $\theta$ is a parameter. Denote $J(x,\theta):=\{i\in[p]:a_{i}(x,\theta)=0\}$. The following conditions are commonly used in parametric optimization for sensitivity analysis [28], although often more relaxed conditions suffice:
  • Mangasarian-Fromovitz Constraint Qualification (MFCQ): Problem (1) satisfies the MFCQ at $(x_{*},\theta_{*})$ if there exists $q\in^{n}$ such that $\frac{\partial a_{i}}{\partial x}(x_{*},\theta_{*})q>0$, for all $i\in J(x_{*},\theta_{*})$.
@@ III Problem Statement
where $f:^{n}\times^{m}\to^{n}$ is the vector field, and $\xi(t)$, $\upsilon(t)$ are the state and control input at time $t$, respectively.
@@
(Unique globally exponentially stable equilibrium): For any constant input signal $\upsilon(t)=u$, system (2) admits a unique equilibrium, which is globally exponentially stable.
This assumption is standard in the literature of feedback optimization and is satisfied in many relevant applications (see, e.g., [1]). It guarantees that controller dynamics can be made slow enough, compared to the plant dynamics, so that stability is not compromised when the controller is interconnected to the plant. For convenience, we let $w:^{m}\to^{n}$ denote the map that assigns each input $u\in^{m}$ to the unique equilibrium $w(u)$, i.e., $f(w(u),u)=0$.
@@
where $\Phi:^{n}\to$, $h:^{n}\to$ and $b:^{m}\to$ are sufficiently smooth functions. The equality constraint $x=w(u)$ is precisely the steady-state constraint. Further, we require the input and state inequality constraints to be respected at all times, that is, for all $t$:
$(\xi(t),\upsilon(t))\in\mathcal{X}:=\{(x,u)\in^{n}\times^{m}:\ h(x)\geq 0,\ b(u)\geq 0\}.$ (4)
@@ IV High-order SGFs for feedback optimization with state constraints
In the following, we denote $\mathcal{L}_{f}h(x,u)=\dfrac{\partial h}{\partial x}(x)\cdot f(x,u)$ and, abusively, $\mathcal{L}^{i}_{f}h(x,u)=\dfrac{\partial\mathcal{L}^{i-1}_{f}h}{\partial x}\cdot f(x,u)$, for $i\geq 2$. We further define $h_{i}:^{n}\times^{m}\to$ for $i\in\mathbb{Z}_{\geq 0}$ as:
$h_{0}(x,u)$ $:=h(x)$ (5)
@@
where $\beta>0$ is a design parameter. In what follows, we assume $\beta>0$ is fixed. We also define the sets $\mathcal{X}_{i}:=\{(x,u)\in^{n}\times^{m}:h(x)\geq 0,b(u)\geq 0,h_{i}(x,u)\geq 0\}$ for $i\in\mathbb{Z}_{\geq 0}$. Note that $\mathcal{X}_{i}\subseteq\mathcal{X}$, for all $i\in\mathbb{Z}_{\geq 0}$, and ${\mathcal{X}}_{0}={\mathcal{X}}$.
@@
$g_{\epsilon,\alpha,\gamma}(x,u)=\left\{\begin{aligned} \operatorname*{arg\,min}_{q}\text{ }&\frac{1}{2}\|q+\epsilon\frac{\partial w}{\partial u}(u)^{\top}\nabla\Phi(x)\|^{2}\\ \mathrm{s.t.:}\text{ }&\nabla^{\top}b(u)\cdot q+\alpha b(u)\geq 0\\ &\begin{aligned} \frac{\partial h_{r}}{\partial x}(x,u)\cdot f(x,u)+\frac{\partial h_{r}}{\partial u}(x,u)\cdot q&\\ +\gamma h_{r}(x,u)\geq 0&\end{aligned}\end{aligned}\right.$ (6)
@@ IV-B Well-posedness: feasibility and Lipschitz continuity
Let Assumption 2 hold. Assume that $\mathcal{X}_{r}$ is compact and that, for any $(x,u)\in\mathcal{X}_{r}$, there exists $q_{(x,u)}\in^{m}$ such that:
@@
(Ensuring CRCQ): It can be shown that CRCQ holds for (6) if Assumption 2 holds, $\mathcal{X}_{r}$ is compact and there exists $c<1$ such that, if $b(u)=0$ and $h_{r}(x,u)=0$, then $|\nabla^{\top}b(u)\frac{\partial h_{r}}{\partial u}(x,u)|\leq c\|\nabla b(u)\|\cdot\|\frac{\partial h_{r}}{\partial u}(x,u)\|$, i.e., $\nabla b(u)$ and $\frac{\partial h_{r}}{\partial u}(x,u)$ are linearly independent in a non-vanishing fashion. The proof is omitted for space limitations.
@@ IV-C Critical points, equilibria and convergence to optimizers
Equivalence between critical points and equilibria on the boundary can be guaranteed if either condition 2 or 3 in Prop. IV.5 holds. These conditions can be interpreted as follows. In the steady state, controller (6) can be seen as the SGF solving the optimization problem $\min_{x,u}\Phi(x)$ s.t. $x=w(u),\ b(u)\geq 0,\ h_{r}(x,u)\geq 0$. Either of these conditions guarantees that this problem is equivalent to the original problem (3). If these conditions are not satisfied, however, the closed-loop dynamics (7) might have equilibria on the boundary which are not critical points. Addressing this is left for future work.
@@
It can be shown that, for any threshold $\delta>0$, there exist $p,\varepsilon>0$ such that the optimizer $(x^{\prime},u^{\prime})$ of (4) satisfies $h(x^{\prime})>0$ and $|\Phi(x^{\prime})-\Phi(x_{*})|<\delta$ (we omit the proof for space limitations). Thus, using the objective function of (4) instead of $\Phi$ in the controller (6), we can invoke Prop. IV.5 to guarantee that the corresponding closed-loop dynamics admit an equilibrium $(x^{\prime},u^{\prime})$ of arbitrarily small suboptimality wrt $(x_{*},u_{*})$. Further, under the assumptions of Thm. IV.6, $(x^{\prime},u^{\prime})$ can be made locally asymptotically stable relative to ${\mathcal{X}}_{r}$.
@@ VII Technical proofs
${\mathcal{N}}_{1}=\bigcup_{(x,u)\in S_{1}}{\mathcal{N}}_{(x,u)},\ {\mathcal{N}}_{2}=\bigcup_{(x,u)\in S_{2}}{\mathcal{N}}_{(x,u)},\ {\mathcal{N}}_{3}=\bigcup_{(x,u)\in S_{3}}{\mathcal{N}}_{(x,u)}.$
@@
    Let $Q_{4}=\mathcal{X}_{r}\setminus(Q_{1}\cup Q_{2}\cup Q_{3})$. Note that for all $(x,u)\in Q_{4}$, we have $b(u)>\min\{c_{1},c_{3}\}>0$ and $h_{r}(x,u)>\min\{c_{2},c_{3}\}>0$.
    Define
@@
By following the argument used in Prop. IV.3, by taking $\alpha>\alpha_{f}$, $\gamma>\gamma_{f}$, (6) is strictly feasible (i.e., for any $(x,u)\in\mathcal{X}_{r}$ there exists $q\in^{m}$ satisfying each of the constraints strictly). Thus, Slater’s condition holds for (6). Since (6) is a strongly convex QP, by [18, Prop. 5.39], we have that MFCQ holds for (6) at the optimizer. The result follows from [28, Thm. 3.6]. ∎
@@
Proof of $\implies$. First, because $x_{*}=w(u_{*})$, we have that $f(x_{*},u_{*})=f(w(u_{*}),u_{*})=0$, from Assumption 2. Due to the assumption on MFCQ or CRCQ, there exist $\lambda_{h},\lambda_{b}\geq 0$ and $\mu\in^{n}$ such that the KKT conditions for (3) are satisfied:
$\begin{bmatrix}\nabla\Phi(x_{*})\\ 0\end{bmatrix}-\lambda_{h}\begin{bmatrix}\nabla h(x_{*})\\ 0\end{bmatrix}-\lambda_{b}\begin{bmatrix}0\\ \nabla b(u_{*})\end{bmatrix}\begin{bmatrix}\mu\\ -\dfrac{\partial w}{\partial u}(u_{*})^{\top}\mu\end{bmatrix}=0,$ (10)
@@
$g_{\epsilon,\alpha,\gamma}(x_{*},u_{*})=\left\{\begin{aligned} &\operatorname*{arg\,min}_{q}\text{ }\frac{1}{2}\|q+\frac{\partial w}{\partial u}(u_{*})^{\top}\nabla\Phi(x_{*})\|^{2}\\ &\quad\mathrm{s.t.:}\text{ }\nabla^{\top}b(u_{*})\cdot q+\alpha b(u_{*})\geq 0\\ &\quad\quad\text{ }\begin{aligned} \underbrace{\frac{\partial h_{r}}{\partial u}(x_{*},u_{*})}_{\frac{\partial\mathcal{L}^{r}_{f}h}{\partial u}(x_{*},u_{*})}\cdot q+\gamma h_{r}(x_{*},u_{*})\geq 0&\end{aligned}\end{aligned}\right.$
@@
where we used (11) and that $h_{r}(x_{*},u_{*})=\beta h_{r-1}(x_{*},u_{*})=\beta^{2}h_{r-2}(x_{*},u_{*})=\dots=\beta^{r}h(x_{*})$. To prove the statement, we need to find $\lambda_{h},\lambda_{b}\geq 0$ and $\mu\in^{n}$ such that (10) holds. We pick $\mu=-\nabla\Phi(x_{*})+\lambda_{h}\nabla h(x_{*})$, and (10) becomes:
@@
First, from [13, Lemma 2.1], under Assumption 1, there exists $W:^{n}\times^{m}\to$, and constants $d_{1},d_{2},d_{3},d_{4}>0$ such that for all $(x,u)\in^{n}\times^{m}$:
@@
Consider the function $V:{\mathcal{N}}\to$ defined by $V(x,u):=\Phi(w(u))-\Phi(w(u_{*}))+W(x,u)$. Let us show that $V$ is a Lyapunov function. Note that $V(x_{*},u_{*})=0$, as $x_{*}=w(u_{*})$, and since $\Phi(w(u))>\Phi(w(u_{*}))$ for all $(x,u)\in{\mathcal{N}}\setminus\{(x_{*},u_{*})\}$, $V(x,u)>0$ for all $(x,u)\in{\mathcal{N}}\setminus\{(x_{*},u_{*})\}$.
@@
where $(\xi(t),\upsilon(t))$ is a solution to (7), with initial condition $(\xi(0),\upsilon(0))\in{\mathcal{X}}_{r}$. Notice that $b(\upsilon(t))\geq 0$ and $h_{r}(\xi(t),\upsilon(t))\geq 0$ for all time, from forward invariance of $\{(x,u):\ b(u)\geq 0,\ h_{r}(x,u)\geq 0\}$ (see proof of Prop. IV.1). In the following, we drop the dependence of $\xi$ and $\upsilon$ on $t$ from notation, for convenience. Let $\Gamma\subseteq{\mathcal{N}}\cap{\mathcal{X}}_{r}$ be a sublevel set of $V$, such that $(x_{*},u_{*})\in\Gamma$ and $h_{r}(x,u)>0$ for all $(x,u)\in\Gamma$ [^4] Such a sublevel set exists because: a) $V$ has compact sublevel sets due to $\Phi$ having compact sublevel sets and $W$ being radially unbounded; and b) $(x_{*},u_{*})$ satisfies $h_{r}(x_{*},u_{*})=\beta^{r}h(x_{*})>0$, and due to continuity of $h_{r}$, $\Gamma$ can be picked small enough so that $h(x)>0$ for all $(x,u)\in\Gamma$. Since $\nabla\Phi$ and $\frac{\partial w}{\partial u}$ are locally Lipschitz, and $\Gamma$ is compact, there exists $L>0$ such that for all $(x,u),(x^{\prime},u)\in\Gamma$, we have $\|\nabla^{\top}\Phi(x)\frac{\partial w}{\partial u}(u)-\nabla^{\top}\Phi(x^{\prime})\frac{\partial w}{\partial u}(u)\|\leq L\|x-x^{\prime}\|$. Then, whenever $(\xi(t),\upsilon(t))\in\Gamma$,
@@
$\zeta(t)=\Big{[}\left\lVert g_{\epsilon,\alpha,\gamma}(\xi(t),\upsilon(t))\right\rVert,\ \left\lVert\xi(t)-w(\upsilon(t))\right\rVert\Big{]},\ P_{\epsilon}=\begin{pmatrix}-\frac{1}{\epsilon}&\frac{(L+d_{4})}{2}\\ \frac{(L+d_{4})}{2}&-d_{3}.\end{pmatrix}$

### v3

@@ I Introduction
  • We prove local asymptotic stability of local optima, assuming that the optimum lies in the interior of the state constraint set;
@@
The existing literature on feedback optimization either disregards state constraints, enforces them only asymptotically, e.g. [4, 10, 11, 12, 13], or considers only static plants, e.g. [14, 15]. In the latter two cases, which are equivalent, state constraints can be enforced through input constraints, as there is a one-to-one correspondence between state and input. Applying these methods to dynamic plants may lead to constraint violations in the transients. In contrast, the method proposed here is the first one to enforce state constraints at all times for dynamic plants. Nonetheless, it comes with local stability guarantees and assumes full knowledge of the plant dynamics, compared to other works (e.g., [9]) that guarantee global convergence to critical points or optimizers, and employ minimal information on system dynamics. Whether our method extends to model-free scenarios and can ensure global convergence remains a topic of future research.
SGFs were first introduced in [16], as dynamical systems that solve constrained optimization problems. These are modified gradient flows, that employ CBFs, to enforce constraint satisfaction. Further, in [19, 20], it was shown how SGFs are continuous approximations of projected gradient flows. SGFs have been used in reinforcement-learning to ensure safety in [21, 22]. In the context of feedback optimization, they have been used as controllers interconnected with dynamic plants in [15, 13], where however state constraints are only asymptotically enforced. Our work is the first to combine SGFs with high-order CBFs. In fact, high-order CBFs are the key to enforce state constraints at all times, since these are of relative degree greater than one with respect to the controller dynamics and cannot be handled by standard SGFs interconnected to a plant, as SGFs (and projected gradient flows) manipulate only the controller’s vector-field. High-order CBFs enable modifications of the controller’s vector field that enforce state constraints.
Finally, while we focus on feedback optimization, there are other methods targeting similar problems, such as extremum seeking (ES; e.g. [23, 24]) and online convex optimization (OCO; e.g. [25, 26, 27, 28]). Like existing feedback-optimization methods, ES methods that address state constraints either cast them as input constraints, implicitly assuming that the plant operates at the steady-state, or consider static plants (which are equivalent approaches); see [23, 24] and references therein. For differences between feedback optimization and ES the reader is referred to [1]. With respect to OCO, the main differences are that OCO generally considers discrete-time plants and focuses on finite-horizon regret guarantees, rather than asymptotic convergence to the optimizer.
@@ II Preliminaries
For $r\in\mathbb{Z}_{>0}$, we denote $r]:=\{1,\ldots,r\}$. We adopt the following conventions on dimensions: given $g:\real^{n}\times\real^{m}\to\real^{k}$, $(x,u)\mapsto g(x,u)$ continuously differentiable, then $\dfrac{\partial g}{\partial x}(x,u)\in\real^{k\times n}$; given $g:\real^{n}\to\real$, $x\mapsto g(x)$ continuously differentiable, then $\nabla g(x)=(\frac{\partial g}{\partial x}(x))^{\top}\in\real^{n}$. Consider the parametric optimization problem
@@
where $\Phi,a_{i}:\real^{n}\times\real^{p}\to\real$ and $\theta$ is a parameter. Denote $J(x,\theta):=\{i\in[p]:a_{i}(x,\theta)=0\}$. The following conditions are commonly used in parametric optimization for sensitivity analysis [29], although often more relaxed conditions suffice:
  • Mangasarian-Fromovitz Constraint Qualification (MFCQ): Problem (1) satisfies the MFCQ at $(x_{*},\theta_{*})$ if there exists $q\in\real^{n}$ such that $\frac{\partial a_{i}}{\partial x}(x_{*},\theta_{*})q>0$, for all $i\in J(x_{*},\theta_{*})$.
@@ III Problem Statement
where $f:\real^{n}\times\real^{m}\to\real^{n}$ is the vector field, and $\xi(t)$, $\upsilon(t)$ are the state and control input at time $t$, respectively.
@@
(Globally exponentially stable equilibrium): For any constant input signal $\upsilon(t)=u$, system (2) admits a unique equilibrium, which is globally exponentially stable.
This assumption is standard in the literature of feedback optimization and is satisfied in many relevant applications (see, e.g., [1]). It guarantees that controller dynamics can be made slow enough, compared to the plant dynamics, so that stability is not compromised when the controller is interconnected to the plant. For convenience, we let $w:\real^{m}\to\real^{n}$ denote the map that assigns each input $u\in\real^{m}$ to the unique equilibrium $w(u)$, i.e., $f(w(u),u)=0$.
@@
where $\Phi:\real^{n}\to\real$, $h:\real^{n}\to\real$ and $b:\real^{m}\to\real$ are sufficiently smooth functions. The equality constraint $x=w(u)$ is precisely the steady-state constraint. Further, we require the input and state inequality constraints to be respected at all times, that is, for all $t$:
$(\xi(t),\upsilon(t))\in\mathcal{X}=\{(x,u)\in\real^{n}\times\real^{m}:\ h(x)\geq 0,\ b(u)\geq 0\}.$ (4)
@@ IV High-order SGFs for feedback optimization with state constraints
In the following, we denote $\mathcal{L}_{f}h(x,u)=\dfrac{\partial h}{\partial x}(x)\cdot f(x,u)$ and, abusively, $\mathcal{L}^{i}_{f}h(x,u)=\dfrac{\partial\mathcal{L}^{i-1}_{f}h}{\partial x}\cdot f(x,u)$, for $i\geq 2$. We further define $h_{i}:\real^{n}\times\real^{m}\to\real$ for $i\in\mathbb{Z}_{\geq 0}$ as:
$h_{0}(x,u)$ $:=h(x),$ (5)
@@
where $\beta>0$ is a design parameter. In what follows, we assume $\beta>0$ is fixed. We also define the sets $\mathcal{X}_{i}:=\{(x,u)\in\real^{n}\times\real^{m}:h(x)\geq 0,b(u)\geq 0,h_{i}(x,u)\geq 0\}$ for $i\in\mathbb{Z}_{\geq 0}$. Note that $\mathcal{X}_{i}\subseteq\mathcal{X}$, for all $i\in\mathbb{Z}_{\geq 0}$, and ${\mathcal{X}}_{0}={\mathcal{X}}$.
@@
$g_{\epsilon,\alpha,\gamma}(x,u)=\left\{\begin{aligned} \operatorname*{arg\,min}_{q}\text{ }&\frac{1}{2}\|q+\epsilon\frac{\partial w}{\partial u}(u)^{\top}\nabla\Phi(x)\|^{2}\\ \mathrm{s.t.:}\text{ }&\nabla^{\top}b(u)\cdot q+\alpha b(u)\geq 0\\ &\begin{aligned} \frac{\partial h_{r}}{\partial x}(x,u)\cdot f(x,u)+\frac{\partial h_{r}}{\partial u}(x,u)\cdot q+\gamma h_{r}(x,u)\geq 0\end{aligned}\end{aligned}\right$. (6)
@@ IV-B Well-posedness: feasibility and Lipschitz continuity
Let Assumption 2 hold. Assume that $\mathcal{X}_{r}$ is compact and that, for any $(x,u)\in\mathcal{X}_{r}$, there exists $q_{(x,u)}\in\real^{m}$ such that:
@@
(Ensuring CRCQ): It can be shown that CRCQ holds for (6) if Assumption 2 holds, $\mathcal{X}_{r}$ is compact and there exists $c<1$ such that, if $b(u)=0$ and $h_{r}(x,u)=0$, then $|\nabla^{\top}b(u)\frac{\partial h_{r}}{\partial u}(x,u)|\leq c\|\nabla b(u)\|\cdot\|\frac{\partial h_{r}}{\partial u}(x,u)\|$, i.e., $\nabla b(u)$ and $\frac{\partial h_{r}}{\partial u}(x,u)$ are linearly independent in a non-vanishing fashion.
@@ IV-C Critical points, equilibria and convergence to optimizers
Equivalence between critical points and equilibria on the boundary is guaranteed if condition 2 or 3 in Prop. IV.5 holds. These conditions can be interpreted as follows. In the steady state, controller (6) can be seen as the SGF solving the optimization problem $\min_{x,u}\Phi(x)$ s.t. $x=w(u),\ b(u)\geq 0,\ h_{r}(x,u)\geq 0$. Either of these conditions enforces that controller (6) is also the SGF of $\min_{u}\Phi(w(u))$ s.t. $b(u)\geq 0,\ h(w(u))\geq 0$, which is the original problem (3). If these conditions are not satisfied, however, the closed-loop dynamics (7) might have equilibria on the boundary which are not critical points. Addressing this is left for future work.
@@
It can be shown that, for any threshold $\delta>0$, there exist $p,\varepsilon>0$ such that the optimizer $(x^{\prime},u^{\prime})$ of (4) satisfies $h(x^{\prime})>0$ and $|\Phi(x^{\prime})-\Phi(x_{*})|<\delta$. Thus, using the objective function of (4) instead of $\Phi$ in the controller (6), we can invoke Prop. IV.5 to guarantee that the corresponding closed-loop dynamics admit an equilibrium $(x^{\prime},u^{\prime})$ of arbitrarily small suboptimality wrt $(x_{*},u_{*})$. Further, under the assumptions of Thm. IV.6, $(x^{\prime},u^{\prime})$ can be made locally asymptotically stable relative to ${\mathcal{X}}_{r}$.
@@ VII Technical proofs
${\mathcal{N}}_{1}=\hskip-5.69054pt\bigcup_{(x,u)\in S_{1}}\hskip-5.69054pt{\mathcal{N}}_{(x,u)},\ {\mathcal{N}}_{2}=\hskip-5.69054pt\bigcup_{(x,u)\in S_{2}}\hskip-5.69054pt{\mathcal{N}}_{(x,u)},\ {\mathcal{N}}_{3}=\hskip-5.69054pt\bigcup_{(x,u)\in S_{3}}\hskip-5.69054pt{\mathcal{N}}_{(x,u)}.$
@@
    Let $Q_{4}=\mathcal{X}_{r}\setminus(Q_{1}\cup Q_{2}\cup Q_{3})$. Note that for all $(x,u)\in Q_{4}$, we have $b(u)>\min\{c_{1},c_{3}\}>0$ and $h_{r}(x,u)>\min\{c_{2},c_{3}\}>0$. Define
@@
By following the argument used in Prop. IV.3, by taking $\alpha>\alpha_{f}$, $\gamma>\gamma_{f}$, (6) is strictly feasible (i.e., for any $(x,u)\in\mathcal{X}_{r}$ there exists $q\in\real^{m}$ satisfying each of the constraints strictly). Thus, Slater’s condition holds for (6). Since (6) is a strongly convex QP, by [18, Prop. 5.39], we have that MFCQ holds for (6) at the optimizer. The result follows from [29, Thm. 3.6]. ∎
@@
Proof of $\implies$. First, because $x_{*}=w(u_{*})$, we have that $f(x_{*},u_{*})=f(w(u_{*}),u_{*})=0$, from Assumption 2. Due to the assumption on MFCQ or CRCQ, there exist $\lambda_{h},\lambda_{b}\geq 0$ and $\mu\in\real^{n}$ such that the KKT conditions for (3) are satisfied:
$\begin{bmatrix}\nabla\Phi(x_{*})\\ 0\end{bmatrix}-\lambda_{h}\begin{bmatrix}\nabla h(x_{*})\\ 0\end{bmatrix}-\lambda_{b}\begin{bmatrix}0\\ \nabla b(u_{*})\end{bmatrix}+\begin{bmatrix}\mu\\ -\dfrac{\partial w}{\partial u}(u_{*})^{\top}\mu\end{bmatrix}=0,$ (10)
@@
[…]
@@
where we used (11) and that $h_{r}(x_{*},u_{*})=\beta h_{r-1}(x_{*},u_{*})=\beta^{2}h_{r-2}(x_{*},u_{*})=\dots=\beta^{r}h(x_{*})$. To prove the statement, we need to find $\lambda_{h},\lambda_{b}\geq 0$ and $\mu\in\real^{n}$ such that (10) holds. We pick $\mu=-\nabla\Phi(x_{*})+\lambda_{h}\nabla h(x_{*})$, and (10) becomes:
@@
First, from [13, Lemma 2.1], under Assumption 1, there exists $W:\real^{n}\times\real^{m}\to\real$, and constants $d_{1},d_{2},d_{3},d_{4}>0$ such that for all $(x,u)\in\real^{n}\times\real^{m}$:
@@
Consider the function $V:{\mathcal{N}}\to\real$ defined by $V(x,u):=\Phi(w(u))-\Phi(w(u_{*}))+W(x,u)$. Let us show that $V$ is a Lyapunov function. Note that $V(x_{*},u_{*})=0$, as $x_{*}=w(u_{*})$, and since $\Phi(w(u))>\Phi(w(u_{*}))$ for all $(x,u)\in{\mathcal{N}}\setminus\{(x_{*},u_{*})\}$, $V(x,u)>0$ for all $(x,u)\in{\mathcal{N}}\setminus\{(x_{*},u_{*})\}$.
@@
where $(\xi(t),\upsilon(t))$ is a solution to (7), with initial condition $(\xi(0),\upsilon(0))\in{\mathcal{X}}_{r}$. Notice that $b(\upsilon(t))\geq 0$ and $h_{r}(\xi(t),\upsilon(t))\geq 0$ for all time, from forward invariance of $\{(x,u):\ b(u)\geq 0,\ h_{r}(x,u)\geq 0\}$ (see proof of Prop. IV.1). In the following, we drop the dependence of $\xi$ and $\upsilon$ on $t$ from notation, for convenience. Let $\Gamma\subseteq{\mathcal{N}}\cap{\mathcal{X}}_{r}$ be a sublevel set of $V$, such that $(x_{*},u_{*})\in\Gamma$ and $h_{r}(x,u)>0$ for all $(x,u)\in\Gamma$. [^4] Such a sublevel set exists because: a) $V$ has compact sublevel sets due to $\Phi$ having compact sublevel sets and $W$ being radially unbounded; and b) $(x_{*},u_{*})$ satisfies $h_{r}(x_{*},u_{*})=\beta^{r}h(x_{*})>0$, and due to continuity of $h_{r}$, $\Gamma$ can be picked small enough so that $h(x)>0$ for all $(x,u)\in\Gamma$. Since $\nabla\Phi$ and $\frac{\partial w}{\partial u}$ are locally Lipschitz, and $\Gamma$ is compact, there exists $L>0$ such that for all $(x,u),(x^{\prime},u)\in\Gamma$, we have $\|\nabla^{\top}\Phi(x)\frac{\partial w}{\partial u}(u)-\nabla^{\top}\Phi(x^{\prime})\frac{\partial w}{\partial u}(u)\|\leq L\|x-x^{\prime}\|$. Then, whenever $(\xi(t),\upsilon(t))\in\Gamma$,
@@
$\zeta(t)=\begin{pmatrix}\left\lVert g_{\epsilon,\alpha,\gamma}(\xi(t),\upsilon(t))\right\rVert&\left\lVert\xi(t)-w(\upsilon(t))\right\rVert\end{pmatrix},\ P_{\epsilon}=\begin{pmatrix}-\frac{1}{\epsilon}&\frac{(L+d_{4})}{2}\\ \frac{(L+d_{4})}{2}&-d_{3}\end{pmatrix}$.